find dy/dx if x x = y y+1 can i do it any three method

 find dy/dx  if xx = yy+1   can i do it any three method 

Grade:12

1 Answers

Vikas TU
6 years ago
Dear student,
Take Log both sides,
xlogx = (y+1)logy
And now differentiate y w.r.t x both sides,
logx + x*1/x = (dy/dx)*logy + (y*1/y)dy/dx + (1/y)dy/dx
1 + logx = (dy/dx)(logy + 1 + 1/y)
dy/dx = (1+logx)/(logy + 1 + 1/y).
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)